OpenStudy (anonymous):

Decompose (x-1)/(3x^2 -14x+15) into partial factors

OpenStudy (luigi0210):

Factor the denominator first

OpenStudy (anonymous):

well I got this for my final answer but my teacher says it wrong... are the denominators factored correctly? \[(3.2\div(3x-5))+(0.5(x-3))\]

OpenStudy (anonymous):

sorry (3.2/(3x-5))+(0.5/(x-3))

OpenStudy (cp9454):

yup thats wrong

OpenStudy (anonymous):

can someone show me how to do it?

OpenStudy (anonymous):

actually i KNOW that my denominators are right but how do I do the numerators

OpenStudy (cp9454):

\[\frac{ x-1 }{ (3x-5)(x-3) } = \frac{ A }{ 3x-5 } + \frac{ B }{x-3 }\]

OpenStudy (cp9454):

solve for A and B, and that will be your answer.

OpenStudy (anonymous):

is this right? 1/(2 (x-3))-1/(2 (3 x-5))

OpenStudy (cp9454):

yes

OpenStudy (anonymous):

k cool

OpenStudy (cp9454):

you got it That's 100% correct